OpenStudy (anonymous):

y≤x-5 y≤-x-4 a. (1, 10) b. (-1, 10) c. (10, 1) d. (1, -10)

OpenStudy (anonymous):

is the question "which one of these points satisfies the inequalities?"

OpenStudy (anonymous):

which of the following points set to the following system of inequalities !

OpenStudy (keenan):

is it a equation or inequalitiy

OpenStudy (anonymous):

if so replace x by the number and y by the number and check. for example if x = 1 and y = 10 the first inequality says \[10 \leq 1-5\] which is false

OpenStudy (anonymous):

inequality

OpenStudy (anonymous):

second one x = -1 and y = 10. first inequality says \[10\leq -1-5\] also false

OpenStudy (anonymous):

third one x = 10, y = 1 first inequality says \[1 \leq 10-5\] which is true. second inequality says \[1 \leq -10-4\] which is false

OpenStudy (anonymous):

finally last one has x = 1, y = -10 first inequality has \[-10\leq 1-5\] true second inequality has \[-10 \leq -1-4\] also true

OpenStudy (anonymous):

so which is the right answer ?

OpenStudy (anonymous):

so only the last ordered pair satisfies both

OpenStudy (anonymous):

hello radar!

OpenStudy (anonymous):

so, D is the answer yes ?

OpenStudy (anonymous):

\[\color{red} {D}\]